A body of mass (m) falls from rest through a medium which exact a frictional drag force (bv) proportional to a applied force (f). 1. Expand the velocity in power series up to t^4


1
Expert's answer
2021-06-07T09:35:18-0400
"ma=m\frac{dv}{dt}=mg-bv+f""\frac{dv}{dt}=g-\frac{b}{m}v+\frac{f}{m}""bdt=\frac{dv}{\frac{mg}{b}+\frac{f}{b}-v}""v(t)=\frac{mg+f}{b}+c\exp{(-bt)}""v(0)=0=\frac{mg+f}{b}+c"

So,


"v(t)=\frac{mg+f}{b}(1-\exp{(-bt)})"

Terminal velocity:


"v(\infty)=\frac{mg+f}{b}"